OpenStudy (haltiamali):

Simplify 2-3(x+1)-4x ..Explain?

OpenStudy (anonymous):

nope jk wrong answer

OpenStudy (anonymous):

Well, you would want to combine all the like terms. This means to get all the variables (x) together, and all the normal numbers together. 1) Use the distributive property to distribute the -3 to the stuff in the parentheses. So: 2-3x-3-4x 2) Now you can combine all the like terms. 2-3= -1 -3x-4x= -7x 3) Final Answer: -1-7x

OpenStudy (haltiamali):

I did it right! :D Thank you! It makes much more sense when you teach..(:

OpenStudy (anonymous):

Good job! You're welcome, lol and thanks :)